rebeccaxhawaii (rebeccaxhawaii):

Derive the equation of the parabola with a focus at (2, -1) and a directrix of y = - one half

rebeccaxhawaii (rebeccaxhawaii):

okay so first ill put it in the formula and take the square root off

rebeccaxhawaii (rebeccaxhawaii):

@haleyelizabeth2017

rebeccaxhawaii (rebeccaxhawaii):

so (x-2)^2+(y+1)^2=(y+1/2)^2

rebeccaxhawaii (rebeccaxhawaii):

right ?

OpenStudy (haleyelizabeth2017):

one sec

OpenStudy (haleyelizabeth2017):

yep

rebeccaxhawaii (rebeccaxhawaii):

okay so then next i distribute y term binomials

OpenStudy (haleyelizabeth2017):

Yes

OpenStudy (haleyelizabeth2017):

What do you get when you simplify everything and isolate y?

rebeccaxhawaii (rebeccaxhawaii):

so i got (x-2)^2=y^2+2y+1=y^2+y+1/4

OpenStudy (haleyelizabeth2017):

you mean a + between the (x-2)^2 and y^2?

rebeccaxhawaii (rebeccaxhawaii):

oops yes sorry

OpenStudy (haleyelizabeth2017):

That is correct.

rebeccaxhawaii (rebeccaxhawaii):

then subract y^2+y+1/4 from both sides

OpenStudy (haleyelizabeth2017):

Nope, subtract the y^2+2y+1 from both sides

rebeccaxhawaii (rebeccaxhawaii):

wait why

OpenStudy (haleyelizabeth2017):

Because you want the (x-2)^2 to be on the opposite side

rebeccaxhawaii (rebeccaxhawaii):

oh okay

rebeccaxhawaii (rebeccaxhawaii):

so its (x-2)^2= -y-3/4

OpenStudy (haleyelizabeth2017):

yes, so now, we can factor out -1 from -y and -3/4. Which gives us\[y+3/4=-(x-2)^2\] but then they will probably have you isolate the y leaving you with \[y=-(x-2)^2-3/4\]

rebeccaxhawaii (rebeccaxhawaii):

okay got it one last one

rebeccaxhawaii (rebeccaxhawaii):

Derive the equation of the parabola with a focus at (-2, 4) and a directrix of y = 6. Put the equation in standard form

rebeccaxhawaii (rebeccaxhawaii):

i didnt get the right answer

rebeccaxhawaii (rebeccaxhawaii):

haleyelizabeth2017

rebeccaxhawaii (rebeccaxhawaii):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

So sorry!

rebeccaxhawaii (rebeccaxhawaii):

lol no its fine

OpenStudy (haleyelizabeth2017):

what did you get when you got this one?

rebeccaxhawaii (rebeccaxhawaii):

well i got 1/20(x+2)^2-1 but thats incorrect

OpenStudy (haleyelizabeth2017):

How'd you get that?

OpenStudy (haleyelizabeth2017):

I got -1/4(x+2)^2+5

rebeccaxhawaii (rebeccaxhawaii):

can you explain one last time how you got that PLEEEEAAAASSSEEE

OpenStudy (haleyelizabeth2017):

\[(x+2)^2+(y-4)^2=(y-6)^2\]\[(x+2)^2+y^2-8y+16=y^2-12y+36\]\[(x+2)^2=-4y+20\]\[(x+2)^2=-4(y-5)\]\[y-5=-1/4(x+2)^2\]\[y=-1/4(x+2)^2+5\]

rebeccaxhawaii (rebeccaxhawaii):

okay thank you so much

OpenStudy (haleyelizabeth2017):

No problem